# DrunkAgent: Stealthy Memory Corruption in LLM-Powered Recommender Agents

### v2

@@
Abstract
Large language model (LLM)-powered agents are increasingly used in recommender systems (RSs) to achieve personalized behavior modeling, where the memory mechanism plays a pivotal role in enabling the agents to autonomously explore, learn and self-evolve from real-world interactions. However, this very mechanism, serving as a contextual repository, inherently exposes an attack surface for potential adversarial manipulations. Despite its central role, the robustness of agentic RSs in the face of such threats remains largely underexplored. Previous works suffer from semantic mismatches or rely on static embeddings or pre-defined prompts, all of which hinder their applicability to systems with dynamic memory states. This challenge is exacerbated by the black-box nature of commercial RSs.
@@ Abstract
1 Introduction
The advent of large language models (LLMs) has ushered a transformative paradigm in recommender systems (RSs). Early LLM-based RSs primarily relied on verbalizing user-item interactions into prompts to guide recommendations (geng2022recommendation, ; bao2023tallrec, ; zhang2023collm, ; wu2024coral, ). However, such approaches often struggle to capture personalized behavioral patterns, due to a fundamental gap between interaction dynamics and generic language modeling (zhang2024agentcf, ). To bridge this gap, recent works have introduced autonomous agents built upon LLMs and augmented with modules, such as memory, planning, and tool-use, laying the foundation for agentic RSs (durante2024agent, ; huang2024foundation, ). As personalized user experiences become central to digital platforms, LLM-powered agentic RSs are rapidly emerging as the next evolutionary step (huang2025towards, ), yet their robustness, particularly against adversarial threats, remains largely underexplored.
At the core of agentic systems lies the memory module, which plays a pivotal role in supporting agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task. More broadly, environment can be any contextual factors that influence the agent’s decisions (zhang2024survey, ). interactions (zhang2024survey, ; wang2024survey, ). Specifically, the module retains historical interactions, allowing agents to accumulate contextual knowledge, adapt to evolving user preferences, and refine their behavior over time through interaction-driven updates (huang2025survey, ), as shown in Fig. 1. However, this very mechanism that empowers agent autonomy also introduces a new and overlooked attack surface.
Adversarial attacks targeting memory often leave a lasting impact. Given that attacks are generally persistent rather than one-off events, they can be readily exploited in non-stationary environments to take advantage of the agent’s continual learning process (huang2025towards, ; zhang2024agentcf, ), leading to repeated integration of adversarial signals into its evolving internal state. This results in long-term memory contamination and hence strengthens a drift in user preferences. Under sequential and retrieval modeling (huang2024foundation, ; huang2025survey, ), these perturbations persist longer by influencing the agent’s ongoing behavior trajectory. Perturbed memory traces can gradually bias the agent’s belief state through temporal dependencies, while also being frequently retrieved in future recommendations. This dual exposure amplifies the longevity and impact of even small-scale attacks. In this work, we take the first step toward enhancing the robustness and security of the LLM-powered agentic RSs by systematically investigating their adversarial vulnerabilities, with a particular focus on stealthy memory corruption attacks.
@@ 1 Introduction
Figure 1: In LLM-powered agentic RSs, agents autonomously interact with their situated environments including other agents to collaboratively optimize their memories over time (Left), with recommendations being generated based on these memories (Right). DrunkAgent perturbs the memory of the target item agent, where the attack strategy aims to ‘get the target agent drunk’ so that the adversarial triggers can be injected into the target memory (The Temporal Snapshot of the Left). As a result, the target item can be promoted to more recommendation lists of normal users (Right).
@@
Existing works on attacks against RSs falls into three largely independent threads: 1) poisoning attacks on traditional collaborative filtering (CF) models (wang2024poisoning, ), which inject fake user profiles to corrupt embeddings; 2) adversarial text attacks on LLM-based recommenders (zhang2024stealthy, ; ning2024cheatagent, ), perturbing input prompts to misguide outputs; and 3) agentic RS architectures that enhance personalization through memory-augmented LLMs (huang2025survey, ). However, these threads remain disconnected. Poisoning methods often assume static embeddings and focus on numerical perturbations, creating a semantic gap with text-driven systems. Textual attacks tend to focus narrowly on pre-defined prompts, limiting their applicability in agentic RSs with dynamic memory updates. Meanwhile, the agentic studies neglect security analyses, leaving critical vulnerabilities unexplored. This challenge is exacerbated by the black-box nature of commercial RSs, where attackers lack access to model parameters and training pipelines. Moreover, a critical threat vector aligned with real-world platform abuse incentives lies in item promotion attacks. Poisoning attacks rely on fake user injections, making promoting items costly and indirect, while textual attacks typically focus on degrading overall performance or causing mis-classification, overlooking practical promotion-driven threats (wang2024poisoning, ; ning2024cheatagent, ; zhang2024stealthy, ).
@@
To tackle the above problems, in this paper, we propose DrunkAgent, the first black-box framework to exploit memory-based vulnerabilities in LLM-powered recommender agents. DrunkAgent operates under two principles: 1) memory confusion, where adversarial textual inputs to disrupt agents’ ability to retain and update interaction histories, leading to persistent distortions in their memories and hence shifts in user preferences, and 2) semantic stealth, where low-perplexity adversarial text are crafted to appear linguistically natural and coherent, enabling undetected memory corruption. Specifically, DrunkAgent crafts semantically meaningful adversarial textual triggers for target item promotions and adopts a series of strategies to corrupt the memory updates of the target item agents during the environmental interaction process, allowing the triggers to achieve maximum impact, as shown in Fig. 1. To mitigate the risk of raising suspicion from frequent queries on victim models in prior methods (zhang2024stealthy, ; ning2024cheatagent, ; wang2024poisoning, ) and to adhere to realistic black-box constraints, DrunkAgent introduces a surrogate model to optimize the triggers and strategies to further improve the attack transferability and stealthiness. To comprehensively evaluate DrunkAgent, we conduct extensive experiments across real-world datasets from different domains, a variety of attack baselines with varying perturbation strengths, representative black-box agentic RSs featuring diverse designs and recommendation tasks, including collaborative filtering, retrieval augmentation, and sequential recommendations, widely adopted stealthiness evaluation methods, and cutting-edge defense mechanisms.
@@
  • We identify a safety vulnerability in autonomous agent-based RSs with powerful LLMs. To the best of our knowledge, we present the first formal and systematic study of adversarial textual attacks on the agentic RSs, aiming to inform future robustness research.
  • We propose DrunkAgent, a novel black-box attack framework specifically designed for the agent-powered RSs, which crafts adversarial triggers and tailored strategies for effective memory perturbations of target item agents.
@@
2 Problem Formulation
2.1 Victim LLM-powered Agentic RSs
@@
We use $\mathcal{Y}=\{y_{u,v}:u\in\mathcal{U},v\in\mathcal{V}\}$ to denote the records of the user-item interaction matrix in the recommendation space, where $\mathcal{U}$ and $\mathcal{V}$ represent the set of real users and the item universe, respectively. $\mathcal{V}_{u}=\{v\in\mathcal{V}:y_{u,v}\neq 0\}$ indicates the set of items that have been interacted by $u$ (i.e., the user’s profile). LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (geng2022recommendation, ; bao2023tallrec, ; zhang2023collm, ; zhao2023recommender, ). Let $\mathcal{X}_{u,v}$ represent the prompts of LLM-powered RSs (denoted as $\text{LLM}_{\Theta}$ with parameters $\Theta$), the preference function can be formulated as below.
$\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$ (1)
@@
where $\mathcal{X}_{u,v}=\mathcal{P}\oplus m_{u}\oplus\mathcal{M}_{v}^{u}\oplus\mathcal{C}_{v}^{u}$ and $\mathcal{R}_{u}$ is the recommendation results for $u$. $\oplus$ denotes the integration of textual strings, encompassing both concatenation and interpolation of substrings. $\mathcal{P}$ stands for prompt templates of recommendations. $m_{u}$ indicates user features and $\mathcal{M}_{v}^{u}=\{m_{v}:v\in\mathcal{V}_{u}\}$ represents the feature set of $u$’s interacted items. $\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (xu2021understanding, ). $|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$\mathcal{K}$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions. In addition to basic textual metadata (e.g., item titles), the features in the vanilla LLM-based RSs include user/item ID information, while the features in the generative agent-based RSs involve the memories of user/item agents. Note that all the elements of $\mathcal{X}_{u,v}$ are optional, which relies on its design.
2.2 Threat Model
2.2.1 Attacker’s Knowledge:
@@
It is often challenging to obtain the internal knowledge of real-world victim RSs, as noted by prior works (ning2024cheatagent, ; yang2024attacking, ; zhang2024stealthy, ). Specifically, critical details such as the parameters (e.g., agents’ memories, which are language-based embeddings (zhang2024agentcf, )) and architectures (e.g., the underlying LLM backbones) are typically inaccessible to adversaries. Given these limitations, DrunkAgent is primarily designed and evaluated under black-box settings (in Section 4). Moreover, due to the inherently open nature of RSs (si2020shilling, ; 10415763, ), in both cases, the attacker is assumed to only have access to publicly available recommendation data such as item titles and user reviews.
2.2.2 Attacker’s Objectives:
@@
As outlined in Section 1, the core objective of our attack is to maximize the exposure of a target item across the widest possible range of normal users. Since black-box attacks are typically optimized in a localized manner (e.g., via surrogate models (wang2024poisoning, )), achieving this objective requires strong transferability, i.e., the ability of the attack to remain effective across different black-box RSs (ning2024cheatagent, ). Furthermore, to broaden the attack impact and maximize the number of affected users in practice, stealthiness becomes a critical secondary objective, ensuring that the attack remains inconspicuous thus difficult to detect (zhang2024stealthy, ).
2.2.3 Attacker’s Capabilities:
@@
In the black-box scenarios, the attacker has no access to the model’s internal states (e.g., user memories and candidate item memories) and can induce memory corruptions by modifying the description of the target item (i.e., the initial memory of the target item agent). Descriptions are often both lengthy and context-rich, which can easily conceal perturbations, as elaborated in Appendix A. Moreover, the profit-driven merchant may frequently update the descriptions to overwrite the target memory that has been optimized over time or construct multiple descriptions of items indicating the same target item. This is feasible, due to the merchants are required to maintain their dedicated APIs (liu2021adversarial, ; yang2024attacking, ; zhang2024stealthy, ).
@@
Figure 2: DrunkAgent Overview. The generation module produces adversarial textual triggers for promoting target items. The strategy module creates adversarial strategies to ‘get the target agents drunk’ to allow the triggers to achieve maximum impact. The triggers and the strategies are optimized on the surrogate module to improve the transferability and stealthiness of black-box attacks.
3 DrunkAgent
@@
3.1 Surrogate Module
This module serves as a surrogate agentic RS (denoted as $f_{\text{Agent}_{\Phi}}(\cdot)$) to evaluate attack effectiveness under limited accessible resources, drawing inspiration from prior works lin2022shilling ; tang2020revisiting ; 10415763 ; yang2024attacking . To avoid frequent queries that could arouse suspicion and to adapt to realistic black-box conditions, the generation module and the strategy module are optimized within this surrogate setting to improve the stealthiness and transferability of DrunkAgent. As discussed in Section 1, agentic RSs are composed of user agents, item agents, and recommendation agents, and all built upon LLM backbones enhanced with memory modules, where the LLMs are used to perform text generation tasks. As such, open-source and auto-regressive LLMs are preferable for constructing surrogate backbones due to their strong generation capabilities and customizability. Given that the most influential Meta’s LLaMA family has been widely adopted in the recent recommendation community (zhao2023recommender, ; bao2023tallrec, ), the latest and powerful LLaMA 3 (dubey2024llama, ) is an ideal choice, which can provide more valuable and instructive feedback for effective optimizations, as illustrated in Fig. 2. The memories for each types of surrogate agents will be detailed in the following sub-sections.
3.2 Generation Module

### v3

@@
Abstract.
Large language model (LLM)-powered agents are increasingly used in recommender systems (RSs) to achieve personalized behavior modeling, where the memory mechanism plays a pivotal role in enabling the agents to autonomously explore, learn and self-evolve from real-world interactions. However, this very mechanism, serving as a contextual repository, inherently exposes an attack surface for potential adversarial manipulations. Despite its central role, the robustness of agentic RSs in the face of such threats remains largely underexplored. Previous works suffer from semantic mismatches or rely on static embeddings or pre-defined prompts, all of which are not designed for dynamic systems, especially for dynamic memory states of LLM agents. This challenge is exacerbated by the black-box nature of commercial recommenders.
@@ Abstract.
Recommender Systems, Adversarial Attacks, Generative Agents, Large Language Models, Collaborative Filtering
copyright: noneccs: Information systems Recommender systemsccs: Security and privacy Web application security
@@
Figure 1. In LLM-powered agentic RSs, agents autonomously interact with their situated environments including other agents to collaboratively optimize their memories over time (Left), with recommendations being generated based on these memories (Right). DrunkAgent perturbs the memory of the target item agent, where the attack strategy aims to ‘get the target agent drunk’ so that the well-designed adversarial triggers can be injected into the target memory (The Temporal Snapshot of the Left). As a result, the target item can be promoted to more recommendation lists of normal users (Right).
1. Introduction
The advent of large language models (LLMs) has ushered a transformative paradigm in recommender systems (RSs). Early LLM-based RSs primarily relied on verbalizing user-item interactions into prompts to guide recommendations (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023; Wu et al., 2024). However, such approaches often struggle to capture personalized behavioral patterns, due to a fundamental gap between interaction dynamics and generic language modeling (Zhang et al., 2024b). To bridge this gap, recent works have introduced autonomous agents built upon LLMs and augmented with modules, such as memory, planning, and tool-use, laying the foundation for agentic RSs (Durante et al., 2024; Huang et al., 2024). As personalized user experiences become central to digital platforms, LLM-powered agentic RSs are rapidly emerging as the next evolutionary step (Huang et al., 2025b), yet their robustness, particularly against adversarial threats, remains largely underexplored.
At the core of agentic systems lies the memory module, which plays a pivotal role in supporting agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task. More broadly, environment can be any contextual factors that influence the agent’s decisions (Zhang et al., 2024a). interactions (Zhang et al., 2024a; Wang et al., 2024b). Specifically, the module retains historical interactions, allowing agents to accumulate contextual knowledge, adapt to evolving user preferences, and refine their behavior over time through interaction-driven updates (Huang et al., 2025a), as shown in Fig. 1. However, this very mechanism that empowers agent autonomy also introduces a new and overlooked attack surface.
Adversarial attacks targeting memory often leave a lasting impact. Given that attacks are generally persistent rather than one-off events, they can be readily exploited in non-stationary environments to take advantage of the agent’s continual learning process (Huang et al., 2025b; Zhang et al., 2024b), leading to repeated integration of adversarial signals into its evolving internal state. This results in long-term memory contamination and hence strengthens a drift in user preferences. Under sequential and retrieval modeling (Huang et al., 2024, 2025a), these perturbations persist longer by influencing the agent’s ongoing behavior trajectory. Perturbed memory traces can gradually bias the agent’s belief state through temporal dependencies, while also being frequently retrieved in future recommendations. This dual exposure amplifies the longevity and impact of even small-scale attacks. In this work, we take the first step toward enhancing the robustness and security of the LLM-powered agentic RSs by systematically investigating their adversarial vulnerabilities, with a particular focus on stealthy memory corruption attacks.
@@ 1. Introduction
Existing works on attacks against RSs falls into three largely independent threads: 1) poisoning attacks on traditional collaborative filtering (CF) models (Wang et al., 2024a), which inject fake user profiles to corrupt embeddings; 2) adversarial text attacks on LLM-based recommenders (Zhang et al., 2024c; Ning et al., 2024), perturbing input prompts to misguide outputs; and 3) agentic RS architectures that enhance personalization through memory-augmented LLMs (Huang et al., 2025a). However, these threads remain disconnected. Poisoning methods often assume static embeddings and focus on numerical perturbations, creating a semantic gap with text-driven systems. Textual attacks tend to focus narrowly on pre-defined prompts, limiting their applicability in agentic RSs, where dynamic memory updates can mitigate the impact of these static attacks to a certain extent. Meanwhile, the agentic studies neglect security analyses, leaving critical vulnerabilities unexplored. This challenge is exacerbated by the black-box nature of commercial RSs, where attackers lack access to model parameters and training pipelines. Moreover, a critical threat vector aligned with real-world platform abuse (e.g., Amazon, eBay, Sony (Lin et al., 2020, 2022)) incentives lies in item promotion attacks. Poisoning attacks rely on fake user injections, making promoting items costly and indirect, while textual attacks typically focus on degrading overall performance or causing mis-classification, overlooking practical promotion-driven threats (Wang et al., 2024a; Ning et al., 2024; Zhang et al., 2024c).
@@
To tackle the above problems, in this paper, we propose DrunkAgent, the first black-box framework to exploit memory-based vulnerabilities in LLM-powered recommender agents. DrunkAgent operates under two principles: 1) memory confusion, where adversarial textual inputs to disrupt agents’ ability to retain and update interaction histories, leading to persistent distortions in their memories and hence shifts in user preferences, and 2) semantic stealth, where low-perplexity adversarial text are crafted to appear linguistically natural and coherent, enabling undetected memory corruption. Specifically, DrunkAgent crafts semantically meaningful adversarial textual triggers for target item promotions and adopts a series of strategies to corrupt the memory updates of the target item agents during the environmental interaction process, allowing the triggers to achieve maximum impact, as shown in Fig. 1. To mitigate the risk of raising suspicion from frequent queries on victim models in prior methods (Zhang et al., 2024c; Ning et al., 2024; Wang et al., 2024a) and to adhere to realistic black-box constraints, DrunkAgent introduces a surrogate model to optimize the triggers and strategies to further improve the attack transferability and stealthiness. To comprehensively evaluate DrunkAgent, we conduct extensive experiments across real-world datasets from different domains, a variety of attack baselines with varying perturbation strengths, representative black-box agentic RSs featuring diverse designs and recommendation tasks, including collaborative filtering, retrieval augmentation, and sequential recommendations, widely adopted stealthiness evaluation methods, and cutting-edge defense mechanisms.
@@
  • We identify a security vulnerability in autonomous, LLM-powered agent-based RSs; specifically, the agent’s memory constitutes a primary attack surface. To the best of our knowledge, we present the first systematic study of adversarial textual attacks on the agentic RSs, aiming to inform future robustness research.
  • We propose DrunkAgent, a novel black-box attack framework tailored for agent-driven RSs. Unlike previous attack methods that only focus on static systems, it effectively perturbs the dynamic memory of the target item agent by crafting adversarial triggers and customized strategies.
@@
2. Problem Formulation
2.1. Victim LLM-powered Agentic RSs
@@
We use $\mathcal{Y}=\{y_{u,v}:u\in\mathcal{U},v\in\mathcal{V}\}$ to denote the records of the user-item interaction matrix in the recommendation space, where $\mathcal{U}$ and $\mathcal{V}$ represent the set of real users and the item universe, respectively. $\mathcal{V}_{u}=\{v\in\mathcal{V}:y_{u,v}\neq 0\}$ indicates the set of items that have been interacted by $u$ (i.e., the user’s profile). LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023; Zhao et al., 2023). Let $\mathcal{X}_{u,v}$ represent the prompts of LLM-powered RSs (denoted as $\text{LLM}_{\Theta}$ with parameters $\Theta$), the preference function can be formulated as below.
(1) $\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$
@@
where $\mathcal{X}_{u,v}=\mathcal{P}\oplus m_{u}\oplus\mathcal{M}_{v}^{u}\oplus\mathcal{C}_{v}^{u}$ and $\mathcal{R}_{u}$ is the recommendation results for $u$. $\oplus$ denotes the integration of textual strings, encompassing both concatenation and interpolation of substrings. $\mathcal{P}$ stands for prompt templates of recommendations. $m_{u}$ indicates user features and $\mathcal{M}_{v}^{u}=\{m_{v}:v\in\mathcal{V}_{u}\}$ represents the feature set of $u$’s interacted items. $\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (Xu et al., 2021). $|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$\mathcal{K}$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions. In addition to basic textual metadata (e.g., item titles), the features in the vanilla LLM-based RSs include user/item ID information, while the features in the generative agent-based RSs involve the memories of user/item agents. Note that all the elements of $\mathcal{X}_{u,v}$ are optional, which relies on its design.
2.2. Threat Model
2.2.1. Attacker’s Knowledge:
@@
It is often challenging to obtain the internal knowledge of real-world victim RSs, as noted by prior works (Ning et al., 2024; Yang et al., 2024; Zhang et al., 2024c). Specifically, critical details such as the parameters (e.g., agents’ memories, which are language-based embeddings (Zhang et al., 2024b)) and architectures (e.g., the underlying LLM backbones) are typically inaccessible to adversaries. Given these limitations, DrunkAgent is primarily designed and evaluated under black-box settings (in Section 4). Moreover, due to the inherently open nature of RSs (Si and Li, 2020; Yang et al., 2023), in both cases, the attacker is assumed to only have access to publicly available recommendation data such as item titles and user reviews.
2.2.2. Attacker’s Objectives:
@@
As outlined in Section 1, the core objective of our attack is to maximize the exposure of a target item across the widest possible range of normal users. Since black-box attacks are typically optimized in a localized manner (e.g., via surrogate models (Wang et al., 2024a)), achieving this objective requires strong transferability, i.e., the ability of the attack to remain effective across different black-box RSs (Ning et al., 2024). Furthermore, to broaden the attack impact and maximize the number of affected users in practice, stealthiness becomes a critical secondary objective, ensuring that the attack remains inconspicuous thus difficult to detect (Zhang et al., 2024c).
2.2.3. Attacker’s Capabilities:
@@
In the black-box scenarios, the attacker has no access to the model’s internal states including user memories and candidate item memories, and can induce memory corruptions by modifying the description of the target item (e.g., the initial memory of the target item agent at $\mathcal{T}_{0}$ in Fig. 1). Descriptions are often both lengthy and context-rich, which can easily conceal perturbations, as elaborated in Section 5. Moreover, the profit-driven merchant may frequently update the descriptions to overwrite the target memory that has been optimized over time or construct multiple descriptions of items indicating the same target item. This is feasible, due to the merchants are required to maintain their dedicated APIs in recommendation platforms (Cohen et al., 2021; Liu and Larson, 2021; Zhang et al., 2024c).
@@
Figure 2. DrunkAgent Overview. The generation module produces adversarial textual triggers for promoting target items. The strategy module creates adversarial strategies to ‘get the target agents drunk’ to allow the triggers to achieve maximum impact. The triggers and the strategies are optimized on the surrogate module to improve the transferability and stealthiness of black-box attacks.
3. DrunkAgent
@@
3.1. Surrogate Module
To evaluate the attack effectiveness under limited accessible resources and to avoid frequent queries that may raise suspicion, we adopt a surrogate module to optimize the black-box attack, which is a commonly used approach for handling realistic black-box constraints (Lin et al., 2022; Tang et al., 2020; Yang et al., 2023, 2024). Within this surrogate setting, we jointly optimize both the generation and strategy modules to enhance the stealthiness and transferability of DrunkAgent. As illustrated in Fig. 2, we construct the module with an agentic RS, denoted as $f_{\text{Agent}_{\Phi}}(\cdot)$, to simulate the behaviors of the real target system. The surrogate framework is composed of user agents, item agents, and recommendation agents, all built upon LLM backbones augmented with memory components. In addition, we incorporate task-aligned prompt templates as behavioral interfaces among the agents, enabling interaction and collaboration. As LLMs are required to perform reasoning and text generation tasks, open-source and auto-regressive LLMs are suitable for constructing the backbones due to their strong generative capabilities and high customizability. The design of memories and templates for different surrogate agents is detailed in the following sub-sections.
3.2. Generation Module